# A Method to Improve the Performance of Reinforcement Learning Based on the $\mathcal{Y}$ Operator for a Class of Stochastic Differential Equation-Based Child-Mother Systems

### v3

@@ 4.1 A Class Of Child-Mother System Modeling
where $O_{t}\in\mathcal{O}\subset\mathbb{R}^{n_{O}}$ is the observation of the whole system and $u_{t}\in\mathcal{U}\subset\mathbb{R}^{d_{m}}$ is the control input of the system. $\mathcal{O}$ is the set of all observations that can be taken into account and $\mathcal{U}$ is the control space of control inputs, i.e., the set of allowed controls. $B^{O}_{t}$ is a $n_{O}$-dimensional Wiener motion. $F(O_{t},u_{t}):\mathbb{R}^{n_{O}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{n_{O}}$, $G(O_{t},u_{t}):\mathbb{R}^{n_{O}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{n_{O}\times n_{O}}$ are the drift term and diffusion term of SDEs, respectively. In this paper, we focus on a class of systems in which the system can be decomposed into a child-mother system as the following form.
$O_{t}=\begin{bmatrix}z_{t}\\ w_{t}\end{bmatrix},\begin{aligned} dz_{t}&=f^{\theta_{z}}(z_{t},u_{t})dt+g^{\theta_{z}}(z_{t},u_{t})dB^{z}_{t},\\ dw_{t}&=\alpha^{\theta_{w}}(w_{t},z_{t})dt+\beta^{\theta_{w}}(w_{t},z_{t})dB^{w}_{t},\\ \end{aligned}$ (39)
where $z_{t}\in\mathcal{Z}\subset\mathbb{R}^{d_{n}}$ is the states of subsystem and $w_{t}\in\mathcal{W}\subset\mathbb{R}^{s_{d}}$ is the states of mother system which means $w_{t}$ will be effected by $z_{t}$. $\mathcal{Z}$ represents the state space of the subsystem state, i.e., the set consisting of all state values that can be taken and $\mathcal{W}$ represents the state space of the state of the mother system, i.e., the set of all values to which the state values of the mother system can be taken. $B^{z}_{t}$ and $B^{w}_{t}$ are the $d_{n}$-dimensional and $s_{d}$-dimensional Wiener motion, respectively. $f^{\theta_{z}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{n}}$ and $g^{\theta_{z}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{n}\times d_{n}}$ are the drift term and diffusion term of SDEs, respectively. Likewise, $\alpha^{\theta_{w}}:\mathbb{R}^{s_{d}}\times\mathbb{R}^{d_{n}}\rightarrow\mathbb{R}^{s_{d}}$ and $\beta^{\theta_{w}}:\mathbb{R}^{s_{d}}\times\mathbb{R}^{d_{n}}\rightarrow\mathbb{R}^{s_{d}\times s_{d}}$ are the drift term and diffusion term of SDEs, respectively. The drift term and diffussion term functions $f^{\theta_{z}},g^{\theta_{z}},\alpha^{\theta_{w}}$ and $\beta^{\theta_{w}}$ of SDEs(39) can be calibrated by the large amount of data observed and collected by the sensors from the system. The parmeters $\theta_{z}$, $\theta_{w}$ are the parameters that the neural network needs to learn. According to Lemma 1, if SDEs(39) has a unique solution, conditions $(i)\sim(iv)$ of Lemma 1 must be obeyed. It’s trivial that conditions $(iii)$ and $(iv)$ are naturally true for the child-mother system in (39). Therefore, the drift term and diffusion term of SDEs in (39) need to satisfy:
@@
$\begin{cases}\mathcal{L}_{L}(\theta_{w})&=\kappa_{1}\int_{0}^{T}(\left\lVert\partial_{w}\alpha^{\theta_{w}}(w_{t},z_{t})\right\rVert_{2}+\left\lVert\partial_{z}\alpha^{\theta_{w}}(w_{t},z_{t})\right\rVert_{2}\\ &+\left\lVert\partial_{w}\beta^{\theta_{w}}(w_{t},z_{t})\right\rVert_{2}+\left\lVert\partial_{z}\beta^{\theta_{w}}(w_{t},z_{t})\right\rVert_{2}-C_{1})^{+}dt,\\ \mathcal{L}_{L}(\theta_{z})&=\kappa_{2}\int_{0}^{T}(\left\lVert\partial_{z}f^{\theta_{z}}(z_{t},u_{t})\right\rVert_{2}+\left\lVert\partial_{u}f^{\theta_{z}}(z_{t},u_{t})\right\rVert_{2}\\ &+\left\lVert\partial_{z}g^{\theta_{z}}(z_{t},u_{t})\right\rVert_{2}+\left\lVert\partial_{u}g^{\theta_{z}}(z_{t},u_{t})\right\rVert_{2}-C_{2})^{+}dt,\\ \end{cases}$ (42)
@@ 4.2 Stochastic Differential Equation Calibration of Child-Mother System
For a typical child-mother system, the deterministic part of the system (in SDEs that is, the drift term function) is, for the most part, uniquely and deterministically described in mathematical form by physical laws. However, the stochastic part of the system is very difficult to be well described, at least under the existing theories, and the uncertainty that these stochastic factors (e.g., complex human behaviors, complex high-dimensional stochastic variables in the environment) bring to the system cannot be described by deterministic dynamic equations (e.g., ordinary differential equations, partial differential equations). Therefore, the modeling of the stochastic part of the system requires a large number of sensors observing the system from reality. And an attempt is made to extract the distribution obeyed by the model of the stochastic part of the system from the large amount of realistic data. In order to make the system discussed in this paper more general, it is assumed that the parameters in the SDEs in (39) are unknown, i.e., the child-mother system is unknown both to the agent in reinforcement learning and to the researchers, and all that is available is the data that the sensors have collected from reality. According to (39), the states of the child-mother system obeys that
@@
where $z_{t+}$, $w_{t+}$ are the states of subsystem and mother system respectively at the next moment of time $t$. It is obvious that systems satisfying SDEs have Markovianity in continuous time. According to Bayesian estimation,
@@
[…]
@@
[…]
@@
[…]
@@
where $\bar{w}_{t}$ is the mother system state values in dataset at time $t$. Therefore, the loss function $\mathcal{L}_{Z}(\theta_{z})$ of the neural network of $f^{\theta_{z}}(\cdot)$ and $g^{\theta_{z}}(\cdot)$ and the loss function $\mathcal{L}_{W}(\theta_{w})$ of the neural network of $\alpha^{\theta_{w}}(\cdot)$ and $\beta^{\theta_{w}}(\cdot)$ are defined as
@@ 5 RL DESIGN BASED ON $\mathcal{Y}$ OPERATOR IN CHILD-MOTHER SYSTEM
Currently, the prevailing reinforcement learning methods are built based on AC framework. The reinforcement learning framework determined by pairing the AC framework with deep neural networks has the advantages that the traditional optimal control lacks, such as in the computation of the value function, the reinforcement learning can be used to estimate the action value function through Critic network, which avoids a large number of computations brought about by the inverse solving of the value function in the traditional optimal control (e.g., dynamic programming), and the difference is obvious when both states and actions are continuous, i.e., the state space and action space are infinite (In mathematics, the potential of the set of state spaces and action spaces is $\aleph_{0}$). Moreover, for systems in continuous state space and continuous action space, in the field of traditional optimal control, it can use the Hamilton-Jacobi-Belman equations for finding the optimal policy, but this method has strong constraints on the design of the system and the cost function. When the model of the controlled system is unknowable or not exhaustible, the traditional optimal control method may be difficult to solve this type of problem effectively. In summary, this chapter proposes a novel reinforcement learning framework based on the $\mathcal{Y}$ operator which mentioned in 3 and applied to the child-mother system built in 4. Moreover, the focus of this paper is not on how to calibrate a system built from SDEs using a large amount of data, but rather on the problem of solving the optimal control for that system once the calibration is complete. The calibration method proposed in 4 is a trivial method, and in the subsequent design of the reinforcement learning framework, by default, the child-mother system has been calibrated by the data, as shown in (50).
@@ 5.1 Critic Network Design
where $\theta_{v}$ is the network parmeter of $V^{\theta_{v}}(z_{t},w_{t})$. The $V^{\theta_{v}}(z_{t},w_{t})$ is the value function of states $z_{t}$ and $w_{t}$ at time $t$, representing the potential value of the state to future total rewards. $R(\cdot)$ is the reward function which is related to the states of child-system $z_{t}$ and mother-system $w_{t}$ as well as the control input $u_{t}$ and the functional form of $R(\cdot)$ should be in the form of a polynomial. $\gamma$ is the reward decay factor due to the fact that rewards at more future moments have a much lower impact on the current action than the current reward has on the action. For the Critic network in the reinforcement learning framework, the task of this network is to give an accurate estimate of the state value function $V^{\theta_{v}}(z_{t},w_{t})$ in the current state so that the Actor network can use the value function estimate $V^{\theta_{v}}(z_{t},w_{t})$ to select the optimal action that maximizes the reward, hence finding a globally optimal solution in the whole control process. According to the theory of reinforcement learning, when the parameter $\theta_{v}$ is trained, the value of $Q(z_{t},u_{t},w_{t})$ is independent of the choice of $s$. It is natural to think of an action $u_{t}$ when it is uniquely determined at time $t$ when the state of the child-system is $z_{t}$ and the state of the mother system is $w_{t}$. Then the action’s value function at that moment should be a deterministic one, i.e., in (51), the value function $Q(z_{t},u_{t},w_{t})$ is independent of the choice of $s$. Therefore, the following equation holds.
@@
00 0 $=\mathbb{E}\left\{R\left(z_{s},u_{s},w_{s}\right)+\ln\gamma V^{\theta_{v}}\left(z_{s},w_{s}\right)\right\}$ (54)
@@
The formula is only related to the reward value at the moment $s$ and the estimation of the value function by the Critic network at the moment $s$, and does not require a partial derivation of the value function, which is a nice property. The method can be remarkably effective when the value function network is not fully understood, such as in some application scenarios of inverse reinforcement learning(IRL), data-driven tasks, etc., where the network structure and parameters of the Critic network are not well known and only a large amount of data is available. According to the theory of $\mathcal{Y}$ operators in Definition 6, the $\mathcal{Y}$ operator with respect to the stochastic process $\{z_{s}\}$ and $V^{\theta_{v}}(z_{s},w_{s})$ can be described by the following equation.
@@
where $z_{s-}$ and $u_{s-}$ are the child-system states and the control input at the last moment of time $s$. It can also be thought of as the historical information available to the entire child-mother system at moment $s$ (i.e., the system history information collected by the sensors). Likewise, the $\mathcal{Y}$ operator with respect to the stochastic process $\{w_{s}\}$ and $V^{\theta_{v}}(z_{s},w_{s})$ can be described by the following equation.
@@
where $w_{s-}$ and $z_{s-}$ are the states of mother-system and the states of child-system states at the last moment of time $s$. The goal of training for the Critic network is essentially to find the optimal parameter $\theta^{*}_{v}$ such that (54) holds at any $s$ moments. Therefore, the loss function of the Critic network can be set intuitively
@@ 5.2 Actor Network Design
where $\mu^{\theta_{u}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{s_{d}}\rightarrow\mathbb{R}^{d_{m}}$, $\sigma^{\theta_{u}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{s_{d}}\rightarrow\mathbb{R}^{d_{m}\times d_{m}}$ are the learning-based function of controller which can be determined by neural network with the parameter $\theta_{u}$. $B^{u}_{\Delta}$ is a $d_{m}$-dimensional random vector. The random variables in random vector $B^{u}_{\Delta}$ are independently and identically distributed random variables, all of which follow the Gaussian distribution with expectation 00 and variance $\Delta T$(time interval of control inputs). Therefore, at time $t$, the action ouputted by Actor network obeys the following distribution:
@@
where $\mu^{\theta_{u}}(z_{t},w_{t})$ is the expectation vector and $\Delta T\cdot\sigma^{\theta_{u}}(z_{t},w_{t})\times\sigma^{\theta_{u}}(z_{t},w_{t})^{T}$ is the covariance matrix. The key to improve the policy of Actor is to find the optimal parmeter $\theta^{*}_{u}$ of Actor network that maximize the reward in reinforcement learning. It means that find
@@
where $\pi^{\theta_{u}}(u_{t}|z_{t},w_{t})$ is the conditional distribution of action $u_{t}$ when the states of child-system $z_{t}$ and states of mother-system $w_{t}$ is well known at time $t$. $\hat{A}_{t}$ is the advantage function, which indicates whether the current action is a favorable action for the current value function, when $\hat{A}_{t}$ is greater than 00, it means that the current action is an advantageous action, on the contrary, the action is a worse action.
According to the Actor design in TRPOschulman2015trust and PPOschulman2017proximal , (65) can be reformulated as
@@
where $\pi^{\theta_{u}}_{old}$ represents the conditional probability of the action under the old strategy. Most current studies comparing the advantages of algorithms use the PPO algorithm as a baseline for comparison. The subsequent experimental results in this paper are also compared with PPO, therefore, the design of Actor takes the same design method as the PPO algorithm. Thus, define $\mathcal{L}_{a}(\theta_{u})$ as the loss function of the Actor as follows
@@
where ”clip” is a function that restricts $\mathcal{R}^{\theta_{u}}_{t}$ to the middle of $1-\epsilon$ and $1+\epsilon$. $\epsilon$ is a hyperparameterization. In practice, during reinforcement learning training, the Actor network and the Critic network are not trained first after the Critic network is trained and then the Actor network is trained. Instead, after the agent interacts with the environment for a fixed number of episodes, the data from these episodes are used, either synchronously or asynchronously, to train the Actor network and Critic network.

### v4

@@ 4.1 A Class Of Child-Mother System Modeling
where $O_{t}\in\mathcal{O}\subset\mathbb{R}^{n_{O}}$ is the observation of the whole system and $u_{t}\in\mathcal{U}\subset\mathbb{R}^{d_{m}}$ is the control input of the system. $\mathcal{O}$ is the set of all observations that can be taken into account and $\mathcal{U}$ is the control space of control inputs, i.e., the set of allowed controls. $B^{O}_{t}$ is a $n_{O}$-dimensional Wiener motion. $F(O_{t},u_{t}):\mathbb{R}^{n_{O}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{n_{O}}$, $G(O_{t},u_{t}):\mathbb{R}^{n_{O}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{n_{O}\times n_{O}}$ are the drift term and diffusion term of SDEs, respectively.
In this paper, we focus on a class of systems in which the system can be decomposed into a child-mother system as the following form.
[…]
where $z_{t}\in\mathcal{Z}\subset\mathbb{R}^{d_{n}}$ is the states of subsystem and $w_{t}\in\mathcal{W}\subset\mathbb{R}^{s_{d}}$ is the states of mother system which means $w_{t}$ will be effected by $z_{t}$. $\mathcal{Z}$ represents the state space of the subsystem state, i.e., the set consisting of all state values that can be taken and $\mathcal{W}$ represents the state space of the state of the mother system, i.e., the set of all values to which the state values of the mother system can be taken. $B^{z}_{t}$ and $B^{w}_{t}$ are the $d_{n}$-dimensional and $s_{d}$-dimensional Wiener motion, respectively. $f^{\theta_{z}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{n}}$ and $g^{\theta_{z}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{n}\times d_{n}}$ are the drift term and diffusion term of SDEs, respectively. Likewise, $\alpha^{\theta_{w}}:\mathbb{R}^{s_{d}}\times\mathbb{R}^{d_{n}}\rightarrow\mathbb{R}^{s_{d}}$ and $\beta^{\theta_{w}}:\mathbb{R}^{s_{d}}\times\mathbb{R}^{d_{n}}\rightarrow\mathbb{R}^{s_{d}\times s_{d}}$ are the drift term and diffusion term of SDEs, respectively.
The drift term and diffussion term functions $f^{\theta_{z}},g^{\theta_{z}},\alpha^{\theta_{w}}$ and $\beta^{\theta_{w}}$ of SDEs(39) can be calibrated by the large amount of data observed and collected by the sensors from the system. The parmeters $\theta_{z}$, $\theta_{w}$ are the parameters that the neural network needs to learn.
According to Lemma 1, if SDEs(39) has a unique solution, conditions $(i)\sim(iv)$ of Lemma 1 must be obeyed. It’s trivial that conditions $(iii)$ and $(iv)$ are naturally true for the child-mother system in (39). Therefore, the drift term and diffusion term of SDEs in (39) need to satisfy:
@@
[…]
@@ 4.2 Stochastic Differential Equation Calibration of Child-Mother System
For a typical child-mother system, the deterministic part of the system (in SDEs that is, the drift term function) is, for the most part, uniquely and deterministically described in mathematical form by physical laws. However, the stochastic part of the system is very difficult to be well described, at least under the existing theories, and the uncertainty that these stochastic factors (e.g., complex human behaviors, complex high-dimensional stochastic variables in the environment) bring to the system cannot be described by deterministic dynamic equations (e.g., ordinary differential equations, partial differential equations).
Therefore, the modeling of the stochastic part of the system requires a large number of sensors observing the system from reality. And an attempt is made to extract the distribution obeyed by the model of the stochastic part of the system from the large amount of realistic data.
In order to make the system discussed in this paper more general, it is assumed that the parameters in the SDEs in (39) are unknown, i.e., the child-mother system is unknown both to the agent in reinforcement learning and to the researchers, and all that is available is the data that the sensors have collected from reality.
According to (39), the states of the child-mother system obeys that
@@
where $z_{t+}$, $w_{t+}$ are the states of subsystem and mother system respectively at the next moment of time $t$. It is obvious that systems satisfying SDEs have Markovianity in continuous time.
According to Bayesian estimation,
@@
$\theta^{*}_{z}$ $=\operatorname*{arg\,max}\limits_{\theta_{z}}\int_{0}^{T}P(\bar{z}_{t+}|\theta_{z})P(\theta_{z})dt,$ (45)
$\theta^{*}_{w}$ $=\operatorname*{arg\,max}\limits_{\theta_{w}}\int_{0}^{T}P(\bar{w}_{t+}|\theta_{w})P(\theta_{w})dt.$
@@
$\mathcal{L}^{Z}_{T}(\theta_{z})=$ $\int_{0}^{T}P(\bar{z}_{t+}|\theta_{z})P(\theta_{z})dt,$ (46)
$=$ $\int_{0}^{T}\frac{e^{I^{Z}_{t}}}{(2\pi)^{\frac{d_{n}}{2}}\lvert\det(g^{\theta_{z}}(\bar{z}_{t},\bar{u}_{t}))\rvert}dt,$
@@
$\mathcal{L}^{W}_{T}(\theta_{w})=$ $\int_{0}^{T}P(\bar{w}_{t}|\theta_{w})P(\theta_{w})dt,$ (47)
$=$ $\int_{0}^{T}\frac{e^{I^{W}_{t}}}{(2\pi)^{\frac{s_{d}}{2}}\lvert\det(\beta^{\theta_{w}}(\bar{w}_{t},\bar{z}_{t}))\rvert}dt,$
@@
where $\bar{w}_{t}$ is the mother system state values in dataset at time $t$.
Therefore, the loss function $\mathcal{L}_{Z}(\theta_{z})$ of the neural network of $f^{\theta_{z}}(\cdot)$ and $g^{\theta_{z}}(\cdot)$ and the loss function $\mathcal{L}_{W}(\theta_{w})$ of the neural network of $\alpha^{\theta_{w}}(\cdot)$ and $\beta^{\theta_{w}}(\cdot)$ are defined as
@@ 5 RL DESIGN BASED ON $\mathcal{Y}$ OPERATOR IN CHILD-MOTHER SYSTEM
Currently, the prevailing reinforcement learning methods are built based on AC framework. The reinforcement learning framework determined by pairing the AC framework with deep neural networks has the advantages that the traditional optimal control lacks, such as in the computation of the value function, the reinforcement learning can be used to estimate the action value function through Critic network, which avoids a large number of computations brought about by the inverse solving of the value function in the traditional optimal control (e.g., dynamic programming), and the difference is obvious when both states and actions are continuous, i.e., the state space and action space are infinite (In mathematics, the potential of the set of state spaces and action spaces is $\aleph_{0}$). Moreover, for systems in continuous state space and continuous action space, in the field of traditional optimal control, it can use the Hamilton-Jacobi-Belman equations for finding the optimal policy, but this method has strong constraints on the design of the system and the cost function. When the model of the controlled system is unknowable or not exhaustible, the traditional optimal control method may be difficult to solve this type of problem effectively.
In summary, this chapter proposes a novel reinforcement learning framework based on the $\mathcal{Y}$ operator which mentioned in 3 and applied to the child-mother system built in 4. Moreover, the focus of this paper is not on how to calibrate a system built from SDEs using a large amount of data, but rather on the problem of solving the optimal control for that system once the calibration is complete. The calibration method proposed in 4 is a trivial method, and in the subsequent design of the reinforcement learning framework, by default, the child-mother system has been calibrated by the data, as shown in (50).
@@ 5.1 Critic Network Design
where $\theta_{v}$ is the network parmeter of $V^{\theta_{v}}(z_{t},w_{t})$. The $V^{\theta_{v}}(z_{t},w_{t})$ is the value function of states $z_{t}$ and $w_{t}$ at time $t$, representing the potential value of the state to future total rewards. $R(\cdot)$ is the reward function which is related to the states of child-system $z_{t}$ and mother-system $w_{t}$ as well as the control input $u_{t}$ and the functional form of $R(\cdot)$ should be in the form of a polynomial. $\gamma$ is the reward decay factor due to the fact that rewards at more future moments have a much lower impact on the current action than the current reward has on the action.
For the Critic network in the reinforcement learning framework, the task of this network is to give an accurate estimate of the state value function $V^{\theta_{v}}(z_{t},w_{t})$ in the current state so that the Actor network can use the value function estimate $V^{\theta_{v}}(z_{t},w_{t})$ to select the optimal action that maximizes the reward, hence finding a globally optimal solution in the whole control process. According to the theory of reinforcement learning, when the parameter $\theta_{v}$ is trained, the value of $Q(z_{t},u_{t},w_{t})$ is independent of the choice of $s$. It is natural to think of an action $u_{t}$ when it is uniquely determined at time $t$ when the state of the child-system is $z_{t}$ and the state of the mother system is $w_{t}$. Then the action’s value function at that moment should be a deterministic one, i.e., in (51), the value function $Q(z_{t},u_{t},w_{t})$ is independent of the choice of $s$. Therefore, the following equation holds.
@@
$0$ $=\mathbb{E}\left\{R\left(z_{s},u_{s},w_{s}\right)+\ln\gamma V^{\theta_{v}}\left(z_{s},w_{s}\right)\right\}$ (54)
@@
The formula is only related to the reward value at the moment $s$ and the estimation of the value function by the Critic network at the moment $s$, and does not require a partial derivation of the value function, which is a nice property. The method can be remarkably effective when the value function network is not fully understood, such as in some application scenarios of inverse reinforcement learning(IRL), data-driven tasks, etc., where the network structure and parameters of the Critic network are not well known and only a large amount of data is available.
According to the theory of $\mathcal{Y}$ operators in Definition 6, the $\mathcal{Y}$ operator with respect to the stochastic process $\{z_{s}\}$ and $V^{\theta_{v}}(z_{s},w_{s})$ can be described by the following equation.
@@
where $z_{s-}$ and $u_{s-}$ are the child-system states and the control input at the last moment of time $s$. It can also be thought of as the historical information available to the entire child-mother system at moment $s$ (i.e., the system history information collected by the sensors).
Likewise, the $\mathcal{Y}$ operator with respect to the stochastic process $\{w_{s}\}$ and $V^{\theta_{v}}(z_{s},w_{s})$ can be described by the following equation.
@@
where $w_{s-}$ and $z_{s-}$ are the states of mother-system and the states of child-system states at the last moment of time $s$.
The goal of training for the Critic network is essentially to find the optimal parameter $\theta^{*}_{v}$ such that (54) holds at any $s$ moments. Therefore, the loss function of the Critic network can be set intuitively
@@ 5.2 Actor Network Design
where $\mu^{\theta_{u}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{s_{d}}\rightarrow\mathbb{R}^{d_{m}}$, $\sigma^{\theta_{u}}:\mathbb{R}^{d_{n}}\times\mathbb{R}^{s_{d}}\rightarrow\mathbb{R}^{d_{m}\times d_{m}}$ are the learning-based function of controller which can be determined by neural network with the parameter $\theta_{u}$. $B^{u}_{\Delta}$ is a $d_{m}$-dimensional random vector. The random variables in random vector $B^{u}_{\Delta}$ are independently and identically distributed random variables, all of which follow the Gaussian distribution with expectation 0 and variance $\Delta T$(time interval of control inputs). Therefore, at time $t$, the action ouputted by Actor network obeys the following distribution:
@@
where $\mu^{\theta_{u}}(z_{t},w_{t})$ is the expectation vector and $\Delta T\cdot\sigma^{\theta_{u}}(z_{t},w_{t})\times\sigma^{\theta_{u}}(z_{t},w_{t})^{T}$ is the covariance matrix.
The key to improve the policy of Actor is to find the optimal parmeter $\theta^{*}_{u}$ of Actor network that maximize the reward in reinforcement learning. It means that find
@@
where $\pi^{\theta_{u}}(u_{t}|z_{t},w_{t})$ is the conditional distribution of action $u_{t}$ when the states of child-system $z_{t}$ and states of mother-system $w_{t}$ is well known at time $t$. $\hat{A}_{t}$ is the advantage function, which indicates whether the current action is a favorable action for the current value function, when $\hat{A}_{t}$ is greater than 0, it means that the current action is an advantageous action, on the contrary, the action is a worse action.
According to the Actor design in TRPO[3] and PPO[5], (65) can be reformulated as
@@
where $\pi^{\theta_{u}}_{old}$ represents the conditional probability of the action under the old strategy.
Most current studies comparing the advantages of algorithms use the PPO algorithm as a baseline for comparison. The subsequent experimental results in this paper are also compared with PPO, therefore, the design of Actor takes the same design method as the PPO algorithm. Thus, define $\mathcal{L}_{a}(\theta_{u})$ as the loss function of the Actor as follows
@@
where ”clip” is a function that restricts $\mathcal{R}^{\theta_{u}}_{t}$ to the middle of $1-\epsilon$ and $1+\epsilon$. $\epsilon$ is a hyperparameterization.
In practice, during reinforcement learning training, the Actor network and the Critic network are not trained first after the Critic network is trained and then the Actor network is trained. Instead, after the agent interacts with the environment for a fixed number of episodes, the data from these episodes are used, either synchronously or asynchronously, to train the Actor network and Critic network.